### v1

# Calibrating the Confidence of Large Language Models by Eliciting Fidelity

Abstract
Large language models optimized with techniques like RLHF have achieved good alignment in being helpful and harmless. However, post-alignment, these language models often exhibit overconfidence, where the expressed confidence does not accurately calibrate with their correctness rate. In this paper, we decompose the language model confidence into the Uncertainty about the question and the Fidelity to the answer generated by language models. Then, we propose a plug-and-play method to estimate the confidence of language models. Our method has shown good calibration performance by conducting experiments with 6 RLHF-LMs on four MCQA datasets. Moreover, we propose two novel metrics, IPR and CE, to evaluate the calibration of the model, and we have conducted a detailed discussion on Truly Well-Calibrated Confidence. Our method could serve as a strong baseline, and we hope that this work will provide some insights into the model confidence calibration.
Calibrating the Confidence of Large Language Models by Eliciting Fidelity
Mozhi Zhang1thanks: Work done during internship at Meituan., Mianqiu Huang1, Rundong Shi2, Linsen Guo 2, Chong Peng 2, Peng Yan 2, Yaqian Zhou1 Xipeng Qiu1thanks: Corresponding author. 1School of Computer Science, Fudan University 2 Meituan {mzzhang22, mqhuang23}@m.fudan.edu.cn {shirundong}@meituan.com {zhouyaqian, xpqiu}@fudan.edu.cn
1 Introduction
Large language models (LLMs) acquire vast world knowledge and demonstrate powerful capabilities through pre-training Brown et al. (2020); OpenAI (2023); Bubeck et al. (2023). With technologies like RLHF Ouyang et al. (2022) and RLAIF Bai et al. (2022); Lee et al. (2023), large language models can become more helpful and harmless to align with human preferences Askell et al. (2021). However, how to build a more honest system has not yet been fully discussed. An honest model should have a certain understanding of the boundary of its knowledge, that is, knowing what it does not know Yin et al. (2023); Yang et al. (2023b). A plausible method is utilizing the calibrated confidence to estimate the knowledge boundary of language models. For pre-trained language models, the per-token logit can already be considered a well-calibrated confidence score, which implies that pre-trained language models (mostly) know what they know Kadavath et al. (2022).
However, recent studies have indicated that language models optimized with techniques like RLHF will exhibit issues of overconfidence Lin et al. (2022a); Kadavath et al. (2022); OpenAI (2023); He et al. (2023); Zhao et al. (2023); Tian et al. (2023); Xiong et al. (2023). This issue could be reflected in Multiple-Choice Question Answering (MCQA) tasks, where the probability of RLHF-LMs generating a token and the likelihood of that token being the correct answer are not well-calibrated. For example, an answer provided by RLHF-LMs with 95% confidence does not mean that there is a 95% probability that the answer is correct. This phenomenon may be due to the optimization objective of RLHF, which is to make the model generate responses aligned with human preferences rather than fitting answers that appear more frequently in the corpus during the pre-training stage.
Figure 1: The reliability diagram of our method. In four different MCQA datasets, our method has demonstrated good calibration effects. The experimental data is derived from GPT-3.5-Turbo.
To alleviate the issue of miscalibration, previous work focuses on two perspectives: the logit-based method and the verbalization-based method. Logit-based methods are usually post-hoc. We need to find a higher temperature (usually above 2.0), known as Temperature-Tuning Guo et al. (2017), to make the distribution of the model’s token logit smoother for mitigating overconfidence Kadavath et al. (2022); He et al. (2023). The verbalization-based method usually requires prompt engineering to elicit the model’s confidence, and it also necessitates the model to have strong Self-Awareness Lin et al. (2022a); Tian et al. (2023); Yin et al. (2023). Aggregating the model’s logit-based and verbalization-based confidence can also calibrate the model confidence to some extent Xiong et al. (2023).
As shown in Figure 2, by replacing the model’s answer with “All other options are wrong.”, we can assess whether the model had high fidelity to its previously given answer. Inspired by this phenomenon, we decompose the language model confidence into two dimensions: the Uncertainty about the question and the Fidelity to the answer generated by language models. First, if the answers provided by language model are consistent under multiple samplings, it indicates that language model has lower uncertainty regarding that question. Thus, we could utilize the information entropy of the frequency distribution of sampled answers to calculate the model’s uncertainty about a question. Second, we design a novel method to estimate the model’s fidelity to each of its sampled answers. Last, the uncertainty regarding question $\mathcal{Q}$ and the fidelity to the answer $a_{i}$ together determine the model’s confidence. As shown in Figure 1, our proposed method achieved good calibration across different MCQA datasets. Meanwhile, our approach does not require knowledge of the model’s per-token log-probability, making it broadly applicable to various commercialized RLHF-LMs, which do not provide the per-token log-probability.
To have a closer look at the calibration of model confidence, we propose two novel metrics for evaluating calibration: 1) Inverse Pair Ratio (IPR), which is the proportion of inverse pairs in the Reliability Diagram. This metric could reflect whether the model is well-calibrated from the perspective of the monotonicity of the Reliability Diagram. If the reliability diagram is monotonic, it indicates that the average accuracy of low-confidence answers is always lower than that of high-confidence answers. 2) Confidence Evenness (CE) refers to the uniformity of the density of each bar in the reliability diagram. Our experimental results indicate that, even within the same dataset, there is a significant difference in the confidence of the answers provided by language models for different questions. This suggests that, after calibration, we could filter out some answers with low confidence to enhance the credibility of the model-generated answers. We summarize our main contributions as follows:
  1. 1)
    Our proposed method could be viewed as a strong baseline for eliciting model confidence, where answer set is known. And the calibrated confidence could be viewed as a soft label.
  2. 2)
    We propose two new metrics, $\mathrm{IPR}$ and $\mathrm{CE}$, to evaluate the calibration of LM’s confidence.
  3. 3)
    We conduct a detailed discussion of a research question: “What kind of Confidence is Truly Well-Calibrated?”, and we hope our discussion can bring some insights to the community.
2 Related Work
Recent work has focused on LLM calibration Lin et al. (2022a); Kadavath et al. (2022); OpenAI (2023). In this section, we will briefly introduce two mainstream methods for eliciting the confidence from language models, namely the Logit-based Method and the Verbalization-based Method.
2.1 Logit-based Method
When we can obtain the per-token logits from language models, we can directly use the probability of generating candidate answers as its confidence.
$\mathrm{Conf}(a_{i})=\frac{\exp(\mathrm{logit}_{a_{i}}/t)}{\sum_{j=1}^{|\mathcal{A}|}\exp(\mathrm{logit}_{a_{j}}/t)},$ (1)
where $t$ is the sampling temperature of language models and $|\mathcal{A}|$ is the size of candidate answer set $\mathcal{A}$. Recent studies indicate that good calibration can be achieved by adjusting the temperature of RLHF-LMs Kadavath et al. (2022); He et al. (2023). However, temperature-scaling Guo et al. (2017) often requires higher temperatures, such as above 2.0 Kadavath et al. (2022), which might cause the outputs of the language models to become too random. When the probabilities for model-generated tokens are inaccessible, a straightforward solution is to deploy sampling and use the frequency of the sampled result to estimate the probability of generating this token. For instance, given a question $\mathcal{Q}$, we could sample $K$ times to acquire a set of answers $\mathcal{A}$ containing $N$ distinct answers, and each answer with an associated frequency $n_{i}$. The probability of the model generating answer $a_{i}$ can be estimated by $\frac{n_{i}}{K}$. Therefore, we could estimate the confidence of language models by $\mathcal{P}_{\mathrm{sampled}}(a_{i})$.
$\mathrm{Conf}(a_{i})=\mathcal{P}_{\mathrm{sampled}}(a_{i})=\frac{n_{i}}{K},a_{i}\in\mathcal{A}$ (2)
Figure 2: If the model’s choice of answer changes after replacing the content of its previous selected option with “All other options are wrong”, it could be considered that the model’s fidelity to its previous answer is not high enough.
2.2 Verbalization-based Method
However, some commercial models, such as ChatGPT and Claude, usually do not provide per-token logits. Benefiting from instruction fine-tuningChung et al. (2022); Zhang et al. (2023), language models could generate responses corresponding to the input instructions. Another intuitive method is to prompt large language models to provide their verbalized confidence along with their responses as follows Jiang et al. (2021); Lin et al. (2022a); Tian et al. (2023):
$(\mathrm{Answer},~{}\mathrm{Conf})=\mathrm{LLM}(\mathrm{Question}),$ (3)
This method requires the model to have a strong ability to follow instructions and strong self-awareness (know whether it knows something or not Yin et al. (2023)). Accordingly, verbalized confidence can be a floating-point number between 0 and 1, i.e., ‘0.8’. And it can be linguistic expressions, i.e., ‘Almost Certain’, ‘About Even’, ‘Unlikely’. Although this method is quite easy to implement, we find various different LMs always tend to output some fixed high confidence expressions.
3 Methodology
In this section, we will introduce the method we propose. Our method does not require any knowledge of the per-token logit of language models or trivial prompt engineering to make the language model output its confidence in a specified format.
Figure 3: Our Proposed Method.
3.1 Sampling
Firstly, as shown in the first step from Figure 3, for question $\mathcal{Q}$, by sampling $K$ times, we can obtain a set of candidate answers $\mathcal{A}$. We take the most frequently occurring answer as the final answer. Meanwhile, we can obtain the frequency distribution $\mathcal{P}_{\mathrm{sampled}}$ of candidate answers.
3.2 Eliciting the Fidelity of Answers
[…]
$\mathbf{Fidelity}_{\mathcal{C}}(a_{i})=\frac{\tau^{i}}{\sum_{i=1}^{|\mathcal{C}|}\tau^{i}},$ (4)
where we usually set $\tau$ as 2. As shown in Figure 3, the answer set $\mathcal{A}$ might include multiple different answers. Consequently, we sequentially replace the candidate answer in $\mathcal{A}$ with “All other options are wrong.” to elicit different hierarchical fidelity chains, as depicted in the second step of Figure 3. The fidelity score of each element $a_{i}$ in every hierarchical fidelity chain $\mathcal{C}_{j}$ can be calculated using (4). Thus, the model’s fidelity of answer $a_{i}$ can be calculated by the weighted average fidelity score across different hierarchical chains. Since the hierarchical fidelity chain is elicited by greedy decoding, the frequency of occurrence of different chains is consistent with the frequency of occurrence of the first element $a_{|\mathcal{C}|}$ from left to right. Therefore, the frequency $\mathcal{P}_{\mathrm{sampled}}(a_{|\mathcal{C}|})$ can be viewed as a proxy for the probability $\mathcal{P}_{\mathrm{sampled}}({\mathcal{C}_{j}})$ of different hierarchical fidelity chains to calculate the overall fidelity score $\mathbf{F}(\cdot)$ of each answer.
$\mathbf{F}(a_{i})=\sum_{j=1}^{|\mathcal{A}|}\mathcal{P}_{\mathrm{sampled}}({\mathcal{C}_{j}})\cdot\mathbf{Fidelity}_{\mathcal{C}_{j}}(a_{i}),$ (5)
3.3 Uncertainty Estimation
As shown in Section 3.1, through sampling, we can obtain the frequency of each answer generated by the model and use it to estimate the generation probability of each answer token. Previous works Kadavath et al. (2022); OpenAI (2023) have revealed that RLHF-LMs often exhibit overconfidence in token generation probability, especially in the temperature range we commonly use, such as between 0 and 1.0. However, these probabilities could still reveal, to some extent, the model’s confidence regarding the current question $\mathcal{Q}$. For instance, if the distribution of $\mathcal{P}_{\mathrm{sampled}}$ is flatter, it indicates that the language model has more significant uncertainty regarding the question $\mathcal{Q}$. An intuitive method is calculating the information entropy of the distribution $\mathcal{P}_{\mathrm{sampled}}$ to estimate the model’s uncertainty about question $\mathcal{Q}$ as follows:
$\mathbf{Uncertainty}(\mathcal{Q})=-\frac{\sum^{\mathrm{M}}_{i=1}p_{i}\cdot\log p_{i}}{\log\mathrm{M}},$ (6)
where $\mathrm{M}$ is the option number of question $\mathcal{Q}$. Since the range of the information entropy for $\mathcal{P}_{\mathrm{sampled}}$ is from 0 to $\log\mathrm{M}$, we normalize the information entropy using $\log\mathrm{M}$.
ARC-Challenge MMLU CommonSenseQA TruthfulQA
Method $\mathrm{ECE}_{10}$downarrow $\mathrm{IPR}_{10}$downarrow $\mathrm{CE}_{10}$uparrow Acc $\uparrow$ $\mathrm{ECE}_{10}$downarrow $\mathrm{IPR}_{10}$downarrow $\mathrm{CE}_{10}$uparrow Acc $\uparrow$ $\mathrm{ECE}_{10}$downarrow $\mathrm{IPR}_{10}$downarrow $\mathrm{CE}_{10}$uparrow Acc $\uparrow$ $\mathrm{ECE}_{10}$downarrow $\mathrm{IPR}_{10}$downarrow $\mathrm{CE}_{10}$uparrow Acc $\uparrow$
GPT-3.5-Turbo
Verb 0.069 0.200 0.681 75.597 0.138 0.200 0.795 59.028 0.087 0.178 0.660 71.253 0.215 0.178 0.792 57.405
Ling 0.083 0.464 0.451 75.683 0.197 0.472 0.441 56.019 0.109 0.250 0.451 71.499 0.271 0.667 0.669 59.241
Sampled 0.095 0.067 0.793 79.266 0.120 0.022 0.922 63.151 0.135 0.067 0.782 74.590 0.147 0.044 0.901 59.333
Ours 0.112 0.139 0.897 79.266 0.088 0.083 0.812 63.151 0.073 0.083 0.812 74.590 0.074 0.133 0.775 59.333
GPT-4-Turbo
Verb 0.080 0.400 0.642 92.833 0.045 0.095 0.706 81.25 0.083 0.111 0.713 83.210 0.056 0.044 0.598 83.109
Ling 0.040 0.036 0.520 89.505 0.066 0.083 0.627 78.762 0.056 0.071 0.637 83.702 0.059 0.139 0.635 79.437
Sampled 0.067 0.200 0.221 92.833 0.153 0.311 0.536 80.324 0.121 0.133 0.541 83.866 0.091 0.178 0.478 87.515
Ours 0.127 0.083 0.757 92.833 0.089 0.083 0.906 80.324 0.109 0.083 0.925 83.866 0.042 0.044 0.764 87.515
Table 1: Experimental results derived from GPT-3.5-Turbo and GPT-4-Turbo. For each column in the table, the closer the color is to blue, the better the calibration. And the closer it is to orange, the worse the performance. We also have bolded the best results, and for the second-best results, we have added an underline beneath them.
ARC-Challenge MMLU CommonSenseQA TruthfulQA
Method $\mathrm{ECE}_{10}$downarrow $\mathrm{IPR}_{10}$downarrow $\mathrm{CE}_{10}$uparrow Acc $\uparrow$ $\mathrm{ECE}_{10}$downarrow $\mathrm{IPR}_{10}$downarrow $\mathrm{CE}_{10}$uparrow Acc $\uparrow$ $\mathrm{ECE}_{10}$downarrow $\mathrm{IPR}_{10}$downarrow $\mathrm{CE}_{10}$uparrow Acc $\uparrow$ $\mathrm{ECE}_{10}$downarrow $\mathrm{IPR}_{10}$downarrow $\mathrm{CE}_{10}$uparrow Acc $\uparrow$
Verb. 0.135 0.178 0.752 58.191 0.199 0.178 0.802 45.891 0.107 0.083 0.806 59.214 0.373 0.133 0.874 26.928
Ling 0.298 0.286 0.613 50.853 0.399 0.333 0.709 30.921 0.097 0.222 0.771 60.770 0.594 0.571 0.681 23.990
Sampled 0.121 0.044 0.890 67.702 0.162 0.067 0.919 52.315 0.110 0.044 0.857 70.762 0.236 0.133 0.891 34.517
Token 0.064 0.067 0.521 67.235 0.135 0.067 0.647 54.803 0.064 0.022 0.477 71.007 0.176 0.133 0.577 34.761
Ours 0.063 0.028 0.887 67.702 0.076 0.028 0.829 52.315 0.051 0.056 0.886 70.762 0.080 0.028 0.704 34.517
Table 2: Experimental results derived from Baichuan2-13B-Chat.
3.4 Confidence Estimation
As demonstrated in Sections 3.2 and 3.3, we can estimate the model’s Uncertainty for a given question $\mathcal{Q}$ and the fidelity $\mathbf{F}(\cdot)$ among its different candidate answers. Therefore, the confidence of the model in its answer $a_{i}$ for question $\mathcal{Q}$ is defined as follows:
$\mathbf{Conf}(\mathcal{Q},a_{i})$ $=\big{(}1-\mathbf{Uncertainty}(\mathcal{Q})\big{)}\cdot\mathbf{F}(a_{i}),$ (7)
4 Experiments
To validate the effectiveness of our proposed method, we conducted experiments on different RLHF-LMs such as GPT-3.5-Turbo [^1] https://openai.com/chatgpt, LLaMA2-Chat Touvron et al. (2023) and Baichuan2-13B-Chat Yang et al. (2023a). To mitigate the influence of the sampling algorithm, unless specifically stated otherwise, we use hyper-parameters with a temperature of 1.0 and set top_p as 1.0.
4.1 Experimental Setting
Dataset.
We have conducted experiments on four MCQA datasets to verify the effectiveness of our proposed confidence estimation method. ARC Clark et al. (2018) is a dataset of 7,787 grade-school-level questions. We use the test split of the ARC-Challenge with 1,172 questions for our experiments. MMLU Hendrycks et al. (2021) is a dataset designed to measure knowledge acquired during pretraining and covers 57 subjects. To reduce the cost of API calls, we sampled $\frac{1}{8}$ of the data for testing for each subject. CommonSenseQA Talmor et al. (2019) is a dataset for commonsense question answering, and we use the validation split with 1,221 questions for experiments. TruthfulQA Lin et al. (2022b) is a dataset that contains 817 questions designed to evaluate language models’ preference to mimic some human falsehoods. All the experiments are conducted under a 00-shot setting.
Metrics.
We utilize multiple metrics to evaluate. We bin the predictions from the model by their confidence and report the $\mathrm{ECE}$ (expected calibration error).
$\mathrm{ECE}_{M}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})|,$ (8)
where $M$ is the bin number of the reliability diagram, $N$ is the overall sample number, and $|B_{m}|$ is the number of samples located in $B_{m}$. This paper also defines two novel metrics to evaluate the calibration. The first one is $\mathrm{IPR}$ (Inverse Pair Ratio), which is used to measure the monotonicity of the reliability diagram. If the reliability diagram is monotonic, it indicates that the average accuracy of answers with low confidence is lower than the average accuracy of answers with high confidence.
$\mathrm{IPR}_{M}=\frac{\mathrm{IP}}{C_{K}^{2}},$ (9)
where $\mathrm{IP}$ is the inverse pair number in the reliable diagram, and $K$ is the bin number with a density larger than 0. Suppose the confidence of language models is always equal to their average accuracy on a specific dataset. In that case, we cannot filter out some answers with low confidence to improve the truthfulness of language models. Therefore, we suggest using the $\mathrm{CE}$ (Confidence Evenness) to evaluate the uniformity of the density of each bar in the reliability diagram.
ARC-Challenge MMLU CommonSenseQA TruthfulQA
Method $\mathrm{ECE}_{10}$downarrow $\mathrm{IPR}_{10}$downarrow $\mathrm{CE}_{10}$uparrow Acc $\uparrow$ $\mathrm{ECE}_{10}$downarrow $\mathrm{IPR}_{10}$downarrow $\mathrm{CE}_{10}$uparrow Acc $\uparrow$ $\mathrm{ECE}_{10}$downarrow $\mathrm{IPR}_{10}$downarrow $\mathrm{CE}_{10}$uparrow Acc $\uparrow$ $\mathrm{ECE}_{10}$downarrow $\mathrm{IPR}_{10}$downarrow $\mathrm{CE}_{10}$uparrow Acc $\uparrow$
LLaMA2-7B-Chat
Verb 0.294 0.083 0.482 45.904 0.325 0.267 0.531 41.551 0.208 0.267 0.516 52.662 0.499 0.200 0.626 21.787
Ling 0.452 0.333 0.283 44.625 0.478 0.357 0.315 38.542 0.385 0.250 0.275 51.597 0.647 0.607 0.406 24.113
Sampled 0.329 0.156 0.781 50.683 0.316 0.222 0.900 43.056 0.294 0.178 0.765 54.627 0.389 0.133 0.875 27.540
Token 0.161 0.156 0.430 50.256 0.224 0.333 0.593 42.419 0.148 0.133 0.417 54.791 0.234 0.289 0.484 27.417
Ours 0.073 0.111 0.921 50.683 0.102 0.167 0.890 43.056 0.053 0.167 0.903 54.627 0.121 0.083 0.762 27.540
LLaMA2-13B-Chat
Verb 0.198 0.143 0.495 57.594 0.286 0.214 0.572 45.614 0.204 0.278 0.497 56.260 0.443 0.167 0.732 27.138
Ling 0.327 0.333 0.393 57.301 0.448 0.333 0.378 45.040 0.316 0.133 0.449 56.692 0.627 0.733 0.508 26.864
Sampled 0.297 0.200 0.653 60.239 0.351 0.267 0.788 47.251 0.287 0.156 0.717 58.722 0.461 0.422 0.798 29.131
Token 0.135 0.178 0.408 59.898 0.225 0.244 0.502 47.512 0.142 0.222 0.403 57.007 0.238 0.200 0.429 30.845
Ours 0.069 0.111 0.886 60.239 0.070 0.083 0.852 47.251 0.043 0.083 0.883 58.722 0.121 0.083 0.762 29.131
LLaMA2-70B-Chat
Verb 0.071 0.286 0.369 70.819 0.236 0.194 0.351 53.183 0.069 0.222 0.286 70.680 0.311 0.028 0.522 43.452
Ling 0.223 0.333 0.119 67.833 0.375 0.333 0.096 51.794 0.189 0.067 0.117 70.106 0.507 0.400 0.289 36.597
Sampled 0.220 0.311 0.475 72.867 0.325 0.289 0.289 56.308 0.212 0.089 0.551 72.809 0.351 0.156 0.622 51.897
Token 0.091 0.200 0.315 73.208 0.190 0.378 0.378 56.597 0.093 0.178 0.339 72.645 0.173 0.267 0.352 52.020
Ours 0.085 0.111 0.908 72.867 0.066 0.083 0.898 56.308 0.094 0.111 0.918 72.809 0.093 0.089 0.804 51.897
Table 3: Experimental results derived from LLaMA-2-Chat.
$\mathrm{CE}_{M}=-\frac{\sum^{\mathrm{M}}_{i=1}p_{i}\cdot\log p_{i}}{\log\mathrm{M}},$ (10)
In this paper, we adopt $10$ equal-size bins to calculate $\mathrm{ECE}_{10}$, $\mathrm{IPR}_{10}$ and $\mathrm{CE}_{10}$. We also report the accuracy on these benchmarks to measure whether calibration reduces the accuracy.
Baselines.
We compared our approach with different baselines for eliciting the confidence of language model. The Verb method involves prompting the model to output a floating-point number between 0 and 1 to represent its confidence immediately after providing an answerTian et al. (2023); Lin et al. (2022a). The Ling method entails having the language model express its confidence level in natural language Tian et al. (2023). Since commercial models like ChatGPT do not provide per-token logits, we employed a sampling technique to estimate the probability of token generation, referred to as the Sampled method. Unless otherwise specified, the Sampled method involves sampling 10 times. For open-source models like LLaMA2-Chat, we directly use the probability of token generation as the measure of the language model’s confidence, which we refer to as the Token method. All the prompt templates we use are shown in Appendix B.
4.2 Main Results
Tables 13 show our experimental results on GPT-3.5-Turbo, GPT-4-Turbo, Baichuan2-13B-Chat, and LLaMA2-Chat. Based on the experimental results, the following conclusions can be drawn:
  1. 1)
    Our proposed method demonstrates a clear improvement over the various baselines in terms of three metrics: $\mathrm{ECE}_{10}$, $\mathrm{IPR}_{10}$, and $\mathrm{CE}_{10}$, which demonstrates the effectiveness of our method.
  2. 2)
    The Verb and Ling methods might, to some extent, impair the language model’s accuracy on multiple-choice question answering tasks, which might be caused by more complicated instructions. Additionally, since the Ling method is more complex, it has a greater impact on the overall accuracy than the Verb method.
  3. 3)
    Similar to the conclusion from Tian et al. (2023), the calibration of the Verb method tends to be better than that of the Ling method. This is because the linguistic expressions used in the Ling method are based on human psychology. However, the confidence represented by the same expression may have a gap between humans and models and among different models and different sentences might mean the same thing Kuhn et al. (2023).
  4. 4)
    The $\mathrm{CE}_{10}$ of the Verbalization-based Method is relatively low, which suggests that language models tends to prefer outputting expressions of certain confidence, such as ‘Highly Likely’, 0.8 and 0.9. This phenomenon can also explain why the $\mathrm{ECE}_{10}$ of the Verbalization-based Method improves when the overall average accuracy of the model is between 70-90%.
Figure 4: Our proposed method achieved well-calibrated results across all temperatures. The experimental results are derived from LLaMA2-13B-Chat. The results from Baichuan2-13B-Chat are presented in Appendix Figure 7.
Figure 5: The experimental results are derived from LLaMA2-Chat.
4.3 Ablation Study
As shown in Table 4, removing $\mathrm{Uncertainty}$ and only relying on $\mathrm{Fidelity}$ to estimate the model’s confidence, we can also achieve comparatively better calibration than other methods. This phenomenon indicates that our proposed method reflects the language model’s Fidelity to its answers very well. Meanwhile, it is difficult to estimate the model’s confidence only depending on $\mathrm{Uncertainty}$. As mentioned in 3.3, $\mathrm{Uncertainty}$ is designed for measuring the model’s uncertainty regarding the question $\mathcal{Q}$, rather than its confidence for a particular answer. In the section 3.2, we utilize (4) to calculate the language model’s normalized fidelity in a hierarchical fidelity chain, where $\tau$ is a hyper-parameter. The larger the value of $\tau$, the lower the estimated fidelity for answers closer to the end of the fidelity chain. Our experiments in Table 4 indicate that setting $\tau$ to around 2 is a relatively appropriate choice for the fidelity estimation process. If $\tau$ is too large, the $\mathrm{ECE}_{10}$ will also increase, which will cause the issue of overconfidence of our estimated confidence.
Method ARC MMLU CSQA TruthfulQA Avg.
Ours 0.069 0.070 0.043 0.121 0.076
w/o. Uncertainty 0.122 0.184 0.115 0.202 0.156
w/o. Fidelity 0.675 0.614 0.704 0.677 0.668
$\tau=1.5$ 0.103 0.064 0.066 0.082 0.079
$\tau=2.0$ (Default) 0.069 0.070 0.043 0.121 0.076
$\tau=2.5$ 0.067 0.089 0.040 0.142 0.085
$\tau=3.0$ 0.074 0.107 0.050 0.155 0.097
$\tau=4.0$ 0.085 0.138 0.075 0.165 0.116
$\tau=5.0$ 0.102 0.158 0.094 0.183 0.134
Best Result (Others) 0.135 0.225 0.142 0.238 0.185
Table 4: Ablation study of our method. The results ($\mathrm{ECE}_{10}$) are derived from LLaMA2-13B-Chat.
5 Analysis and Discussion
To take a closer look at the difference between different calibration methods tailored for language models, in this section, we verify the robustness of our method from two aspects: Temperature-Scaling and Parameter-Scaling. Meanwhile, we also conducted a detailed discussion of a research question: What kind of Confidence is Truly Well-Calibrated?
Figure 6: Reliability diagrams of Baichuan2-13B-Chat on ARC-Challenge. In these diagrams, the darker the color, the higher the density. The reliability diagrams of other models we evaluated are shown in Appendix Figures 813.
5.1 Temperature-Scaling
In the main experiments, we evaluate various methods using a constant temperature of 1.0. In this section, we will explore the influence of sampling temperature on the performance of different methods. As illustrated in Figures 4 and 7, our proposed calibration method consistently achieves the lowest expected calibration error across all temperatures, showing remarkable robustness to temperature variations. This is because, in eliciting model fidelity, our method always employs Greedy Decoding rather than Sampling. Thus, the hierarchical chains we obtain are usually consistent across different sampling temperatures. In contrast, the expected calibration error of Logit-based Methods is usually affected by temperature. For the Sampling method with limited sampling budgets, the lower the temperature, the more significantly the diversity of the sampled results will decrease, exacerbating the overconfidence of language models. For the Token Method, the impact of temperature on its calibration shows a trend of “first increasing and then remaining relatively stable” or ”first increasing and then decreasing“. This is because we could directly utilize (1) to estimate the confidence of each option, and if the temperature is too low (i.e., 0.1), it will lead to the confidence of a large number of options approaching zero. This phenomenon might contribute to reducing expected calibration error, but it does not necessarily indicate that the model’s confidence is well-calibrated. The Verbalization-based method is less affected by temperature, which indicates that the expressions which language models prefer to output are relatively consistent across different temperatures.
5.2 Parameter-Scaling
As shown in Figure 5, we evaluate the calibration of various methods at different parameter scales on the LLaMA2-Chat series models. Our proposed method exhibits good calibration across different amounts of model parameters. With the size of model parameters increasing, the calibration of the Verbalization-based method and the Logit-based method is improving. This phenomenon indicates that as the scale of model parameters increases, the model’s Self-Awareness is improving. However, the relatively high expected calibration error suggests that language models still have issues with overconfidence.
5.3 Truly Well-Calibrated Confidence
Previous work mainly evaluates the calibration of language models through expected calibration error (ECE). This section will discuss the research question: “What Kind of Confidence is Truly Well-Calibrated?”. Figure 6 demonstrates the calibration of various methods. From the calibration perspective, we hope that the confidence and accuracy relationship is close to the curve $y=x$. Thus, we need to reduce the ECE by calibrating confidence. Meanwhile, we hope that the reliability diagram should be as monotonic as possible to ensure that the accuracy of the results generated with low confidence is lower than that of the results with high confidence. Therefore, we propose the Inverse Pair Ratio (IPR) to evaluate monotonicity. From the perspective of building a more honest system, we hope the model’s confidence should be more evenly distributed across different confidence intervals. For example, if a language model has an overall accuracy of 75% on the TruthfulQA dataset and the confidence of each question from the language model is always 75%, its ECE and IPR would be 0. In this case, we think that the confidence may not necessarily be a truly well-calibrated confidence because we could not exclude some low-confidence results based on the confidence from the language model. Although the prior distribution of the model’s confidence is unknown, our confidence estimation method finds that language models have different confidence for different questions and different answers to the same question. Thus, we propose a metric called Confidence Evenness (CE) to measure the distance between the posterior confidence distribution and the uniform distribution. We believe ECE, IPR, and CE evaluate calibration from different perspectives and there is a trade-off between these three metrics. We suggest that truly well-calibrated confidence should achieve a balance among these three metrics, ECE, IPR, and CE, rather than over-optimizing any of them.
6 Conclusion
In this paper, we decompose the language model confidence into the Uncertainty about the question and the Fidelity to the answer generated by language models. Through the decomposition, we propose a plug-and-play method to calibrate the confidence of language models. Through experiments with 6 RLHF-LMs on 4 multiple-choice question answering benchmarks, our method exhibits good calibration. Besides, we propose two novel metrics, $\mathrm{IPR}$ and $\mathrm{CE}$, to evaluate the calibration of language models. Finally, we conduct a detailed discussion on Truly Well-Calibrated Confidence. We believe our method can serve as a strong baseline, and we hope that this work will provide some insights into the language model confidence calibration.
Limitations
Although our method has shown good calibration, it is mainly applicable to scenarios where the set of answers is known, i.e., multiple-choice question answering, text classification, sentiment classification, and preference labeling in RLHF. Eliciting the model’s fidelity in open-ended generation scenarios is a direction worth exploring. Meanwhile, our method involves multiple invocations of language models, and how to estimate the probability distribution of tokens generated by the language model with as few callings as possible remains to be studied.
Appendix A Algorithm
The pseudo code of our proposed method is shown in Algorithm 1. It should be clarified that, as long as a candidate answer $a_{i}$ appears in the answer set $\mathcal{A}$ or the Fidelity chain set $\mathcal{S}$, we could estimate its confidence through (7).
Algorithm 1 Algorithm
1:Input question $\mathcal{Q}$, Option list $\mathcal{O}$, Answer set $\mathcal{A}=\varnothing$, Sampling budget $K$, RLHF-LM $\mathrm{LM}$, $o^{*}$ is “All other options are wrong.”, Fidelity chain set $\mathcal{S}$, $\mathbf{U}(\cdot)$ refers to (6).
2:$t\leftarrow 0$
3:while $t<K$ do
4:     $a_{i}\leftarrow\mathrm{LM}(Q,\mathcal{O})$triangleright Sampling answer
5:     $\mathcal{A}\leftarrow\mathcal{A}\cup\{a_{i}\}$
6:     $\mathcal{P}_{sampled}(a_{i})\leftarrow\mathcal{P}_{sampled}(a_{i})+1$
7:     $t\leftarrow t+1$triangleright Continue sampling
8:end while
9:$\mathcal{P}_{sampled}(a_{i})\leftarrow\mathcal{P}_{sampled}(a_{i})/K$
10:
11:$\mathbf{Uncertainty}(\mathcal{Q})=\mathbf{U}(\mathcal{P}_{sampled})$triangleright Get uncertainty
12:$i\leftarrow 0$
13:while $|\mathcal{A}|>0$ do
14:     $\mathcal{A}\leftarrow A\setminus\{a_{i}\}$triangleright Select a answer
15:     $\mathcal{O}^{*}\leftarrow(\mathcal{O}\setminus\{o_{i}\})\cup o_{*}$triangleright Replace option
16:     $\mathcal{C}_{i}=a_{i}$triangleright Init a fidelity chain
17:     while $|\mathcal{O}^{*}|>0$ do
18:         $a^{*}\leftarrow\mathrm{LM}(\mathcal{Q},\mathcal{O}^{*})$triangleright Greedy decoding
19:         if $a^{*}\neq a_{i}$ then $\triangleright$ Low fidelity
20:              $\mathcal{O}^{*}\leftarrow\mathcal{O}^{*}\setminus\{o_{i}\}$triangleright Delete option
21:              $a_{i}=a^{*}$
22:              $\mathcal{C}_{i}=(\mathcal{C}_{i}\rightarrow a_{*})$triangleright Add element
23:         else
24:              break $\triangleright$ High fidelity
25:         end if
26:     end while
27:     $\mathcal{S}\leftarrow\mathcal{S}\cup\mathcal{C}_{i}$
28:     $i\leftarrow i+1$
29:end while
30:
31:$\mathbf{F}(a_{i})=\sum_{j=1}^{|\mathcal{A}|}\mathcal{P}_{\mathrm{sampled}}({\mathcal{C}_{j}})\cdot\mathbf{Fidelity}_{\mathcal{C}_{j}}(a_{i})$triangleright Get fidelity
32:$\mathrm{Conf}(\mathcal{Q},a_{i})=(1-\mathbf{Uncertainty}(\mathcal{Q}))\cdot\mathbf{F}(a_{i})$triangleright Get confidence
33:return $\mathrm{Conf}(\mathcal{Q},a_{i})$triangleright Return the confidence of answer $a_{i}$
Appendix B Prompt Templates
We use the prompt template from  Tian et al. (2023) for a fair comparison. The prompt template for each baseline is provided in Table 7. The question is substituted for the variable \${THE_QUESTION} in each prompt. Table 5 shows the linguistic expression list of confidence we used for the Ling Method, which originates from Fagen-Ulmschneider (2023).
Linguistic Expression Confidence Score
‘Certain’ 1.0
‘Almost Certain’ 0.95
‘Highly Likely’ 0.9
‘Very Good Chance’ 0.8
‘We Believe’ 0.75
‘Probably’ 0.7
‘Probable’ 0.7
‘Likely’ 0.7
‘Better than Even’ 0.6
‘About Even’ 0.5
‘Probably Not’ 0.25
‘We Doubt’ 0.2
‘Unlikely’ 0.2
‘Little Chance’ 0.1
‘Chances are Slight’ 0.1
‘Improbable’ 0.1
‘Highly Unlikely’ 0.05
‘Almost No Chance’ 0.02
‘Impossible’ 0.0
Table 5: The EXPRESSION_LIST we used for the Ling Method.
Figure 7: The Impact of Temperature on Different Methods. Our proposed method achieved well-calibrated results across all temperatures. The experimental results are derived from Baichuan2-13B-Chat.
Appendix C Why could CE be used as a metric?
As mentioned in section  4.2, we found that Language models tend to prefer outputting expressions of certain confidence, such as ’Highly Likely’, 0.8, and 0.9. In the table 8, we have counted the occurrence of different confidence levels for various models on different datasets to demonstrate the model’s preference for certain confidence levels when using the Verb and Ling method.
We also notice that as the model parameters increased, the accuracy of the model improved, but the language model’s preference for certain confidence levels do not change and even became stronger. Therefore, we introduced the Confidence Evenness to assess whether the model’s confidence is overly concentrated in certain intervals.
Can existing metrics (such as ECE) capture this phenomenon? There is an example: on CommonSenseQA, as the parameters of Llama2-Chat increasing, the accuracy rises from 51% to 70%, and the ECE using the Ling method decrease from 0.385 to 0.189. But the 70B model shows a stronger preference for outputting a confidence of 0.9. Focusing solely on the ECE metric cannot fully observe the changes in model preferences. Fortunately, this phenomenal could be reflected by the CE metrics.
Another extreme case is if models of varying parameter sizes always output a 0.9 confidence level, and as the model size increases, the average accuracy just shifts from 70% to 90%, then the ECE would drop to 0. If we only use existing metrics for observation, we might conclude that the model with the largest parameters has the strongest self-awareness. However, by evaluating the CE metric across different models, we can identify a potential preference in how models express confidence. Its ECE becoming 0 might just coincidentally be because the average accuracy on a certain dataset equals the confidence level it prefers to output. Therefore, we believe the CE metric provides a new perspective for observing model confidence calibration.
Finally, it should be noted that we believe an over-concentration of model confidence in a particular value or interval is not conducive to using model confidence as a simple metric to filter out low-confidence answers.
Appendix D The Computation Cost of Eliciting Fidelity
In this section, we will display the average length of the fidelity chains for different models across various datasets in the Table 6. Since we deploy greedy decoding during the process of eliciting fidelity,the average length of the fidelity chain is equal to the average number of requests. At the same time, it should be noted that, when eliciting the Fidelity Chain, only 1 token needs to be generated. Therefore, the average length of the fidelity chain can also be regarded as the average number of tokens generated.
Model ARC-Challenge MMLU CommonSenseQA TruthfulQA Avg.
GPT-3.5-Turbo 2.774 2.984 3.052 3.275 3.021
GPT-4-Turbo 1.492 1.915 2.157 1.616 1.795
Baichuan2-13B-Chat 2.830 2.820 2.889 4.345 3.221
Llama2-7B-Chat 2.467 2.631 2.771 3.944 2.953
Llama2-13B-Chat 2.725 2.875 2.956 4.100 3.164
Llama2-70B-Chat 2.384 2.563 2.455 3.284 2.671
Table 6: The average length of the fidelity chains for different models across various datasets
Appendix E Reliability Diagram
We provide the reliability diagrams of all the RLHF-LMs we evaluated in Figures 8-13. In a reliability diagram, the darker the color of the bar, the greater its density is, which indicates a preference for the confidence the language models express. Although the average accuracy of various RLHF-LMs is quite different, these model always prefer to express their confidence about 70-90% in verbalized methods.
Appendix F Additional Results
Besides the ECE metric, the Brier Score is also commonly used as an evaluation criterion for model calibration.
$\mathrm{BrierScore}_{M}=\frac{1}{N}\sum_{t=1}^{N}(f_{t}-o_{t})^{2},$ (11)
where $f_{t}$ is the probability and $o_{t}$ is the label. Accordingly, $f_{t}$ can be referred to as the model’s confidence, while $o_{t}$ represents whether it is the correct answer (0 indicating an incorrect answer, 1 indicating a correct answer). In Table 9, we present the Brier Scores of various baselines and our proposed method. It can be seen that our method still exhibits good calibration, especially for closed-source models such as GPT-3.5-Turbo, GPT-4 Turbo.
Method Prompt Template
[…]
Table 7: Prompt templates for each method evaluated.
Dataset Method Model $0.0$ $0.02$ $0.05$ $0.1$ $0.2$ $0.25$ $0.3$ $0.4$ $0.5$ $0.6$ $0.7$ $0.8$ $0.9$ $0.95$ $1.0$ $\mathrm{ECE}_{10}$downarrow $\mathrm{CE}_{10}$uparrow Acc $\uparrow$
CSQA Verb LLaMA2-7B-Chat 3 0 0 1 25 0 23 5 78 10 309 727 19 0 21 0.208 0.516 52.662
LLaMA2-13B-Chat 11 0 0 0 9 0 1 29 7 112 108 851 61 0 32 0.204 0.497 56.260
LLaMA2-70B-Chat 6 0 0 2 2 0 3 3 1 23 221 955 2 0 3 0.069 0.286 70.680
Ling LLaMA2-7B-Chat 11 0 21 0 3 0 0 0 1 5 2 13 1020 75 70 0.385 0.275 51.597
LLaMA2-13B-Chat 18 1 11 0 6 0 0 0 0 0 3 194 892 96 0 0.316 0.449 56.692
LLaMA2-70B-Chat 0 0 26 0 0 0 0 0 0 1 2 2 1172 2 16 0.189 0.117 70.106
MMLU Verb LLaMA2-7B-Chat 14 0 0 3 46 0 21 16 65 44 488 981 26 0 24 0.325 0.531 41.551
LLaMA2-13B-Chat 23 0 0 0 41 0 0 54 7 227 278 1056 18 0 24 0.286 0.572 45.614
LLaMA2-70B-Chat 1 0 0 0 7 0 3 1 2 9 518 1159 1 0 27 0.236 0.351 53.183
Ling LLaMA2-7B-Chat 47 0 101 0 21 0 0 0 6 4 7 12 1408 77 45 0.478 0.315 38.542
LLaMA2-13B-Chat 81 1 15 0 4 2 0 0 0 0 4 84 1261 261 11 0.448 0.378 45.040
LLaMA2-70B-Chat 3 0 31 0 0 0 0 0 0 6 2 5 1673 1 7 0.375 0.096 51.794
ARC Verb LLaMA2-7B-Chat 4 0 0 0 26 0 13 6 53 5 216 800 20 0 29 0.294 0.482 45.904
LLaMA2-13B-Chat 1 0 0 0 31 0 0 13 13 68 129 851 18 0 47 0.198 0.495 57.594
LLaMA2-70B-Chat 3 0 0 0 11 0 3 0 2 6 288 836 3 0 20 0.071 0.369 70.819
Ling LLaMA2-7B-Chat 3 0 24 0 10 0 0 0 0 0 5 10 1023 53 44 0.452 0.283 44.625
LLaMA2-13B-Chat 1 0 5 0 5 0 0 0 0 0 1 76 914 162 8 0.327 0.393 57.301
LLaMA2-70B-Chat 3 0 27 1 0 0 0 0 0 3 1 1 1121 2 13 0.223 0.119 67.833
TruthfulQA Verb LLaMA2-7B-Chat 10 0 0 1 23 0 8 2 125 18 167 406 17 0 40 0.499 0.626 21.787
LLaMA2-13B-Chat 11 0 0 1 11 0 0 56 34 145 116 369 26 0 48 0.443 0.732 27.138
LLaMA2-70B-Chat 3 0 0 0 7 0 4 4 4 22 320 404 9 0 30 0.311 0.522 43.452
Ling LLaMA2-7B-Chat 30 0 53 0 10 0 0 0 8 4 4 15 611 43 39 0.647 0.406 24.113
LLaMA2-13B-Chat 39 2 19 0 4 0 0 0 0 0 4 40 526 177 6 0.627 0.508 26.864
LLaMA2-70B-Chat 10 0 31 0 0 0 0 0 0 3 0 9 718 12 31 0.507 0.289 36.597
Table 8: Language models tend to prefer outputting expressions of certain confidence, such as 0.8, and 0.9.
Model Method ARC-Challenge MMLU CommonSenseQA TruthfulQA Avg.
GPT-3.5-Turbo Verb 0.181 0.247 0.189 0.274 0.223
Ling 0.197 0.278 0.204 0.318 0.249
Sampled 0.157 0.202 0.216 0.206 0.195
Ours 0.141 0.170 0.139 0.153 0.151
GPT-4-Turbo Verb 0.181 0.247 0.204 0.274 0.227
Ling 0.198 0.278 0.216 0.318 0.253
Sampled 0.074 0.174 0.147 0.112 0.127
Ours 0.095 0.142 0.134 0.102 0.118
Baichuan2-13B-Chat Verb 0.257 0.294 0.239 0.363 0.288
Ling 0.336 0.407 0.235 0.553 0.383
Sampled 0.196 0.236 0.186 0.262 0.220
Token 0.095 0.168 0.092 0.198 0.138
Ours 0.166 0.193 0.153 0.149 0.165
Llama2-7B-Chat Verb 0.332 0.348 0.283 0.449 0.353
Ling 0.451 0.471 0.396 0.609 0.4821
Sampled 0.358 0.350 0.323 0.411 0.360
Token 0.171 0.238 0.158 0.246 0.203
Ours 0.204 0.214 0.181 0.186 0.196
Llama2-13B-Chat Verb 0.277 0.320 0.272 0.394 0.316
Ling 0.352 0.448 0.343 0.599 0.435
Sampled 0.318 0.374 0.317 0.470 0.370
Token 0.141 0.233 0.150 0.242 0.192
Ours 0.178 0.196 0.166 0.180 0.180
Llama2-70B-Chat Verb 0.206 0.297 0.208 0.332 0.261
Ling 0.267 0.390 0.240 0.496 0.348
Sampled 0.236 0.347 0.237 0.360 0.295
Token 0.094 0.196 0.098 0.174 0.141
Ours 0.154 0.189 0.156 0.162 0.165
Table 9: The Brier Score of different methods from six RLHF-Models on four MCQA datasets.
Figure 8: The experimental results are derived from GPT-3.5-Turbo on 4 MCQA datasets.
Figure 9: The experimental results are derived from GPT-4-Turbo on 4 MCQA datasets.
Figure 10: The experimental results are derived from Baichuan2-13B-Chat on 4 MCQA datasets.
Figure 11: The experimental results are derived from LLaMA2-7B-Chat on 4 MCQA datasets.
Figure 12: The experimental results are derived from LLaMA2-13B-Chat on 4 MCQA datasets.
Figure 13: The experimental results are derived from LLaMA2-70B-Chat on 4 MCQA datasets.